A particle performs a simple harmonic motion about a point x=0m, with amplitude 0.5m and frequency 4Hz. Given that it is momentarily at rest at x=0.5m with t=0s. Determine, the time required for the particle to move from point x=0 to a point 0.3m away.
1
Expert's answer
2020-06-26T14:37:46-0400

"y=A\cos(\omega t+\phi)"


"t=0 \to y=0.5"


So,


"\phi=0"


"y=0.5\cos(2\pi\nu t)=0.5\cos(8\pi t)"


If "y=0" then "t_0=1\/16s=0.037s"


If "y=0.3" then "t=\frac{1}{8\pi}\cos^{-1}(0.6)=0.0625s"


Finally


"\Delta t=0.0625-0.037=0.0255s"